OpenStudy (anonymous):

In a certain advisory there are nine fifteen year olds, seven sixteen year olds and two seventeen year olds. How old is the average advisory member (rounded to the nearest tenth)?

OpenStudy (anonymous):

15.6

Parth (parthkohli):

Can you plot a table with the age, frequency and \(\rm age \times frequency\)?

OpenStudy (anonymous):

((9*15)+(7*16)+(2*17))/(9+7+2) = average age (135+112+34)/(18) = average age 281/81 = 15.6 years old

Parth (parthkohli):

Round to the nearest tenth.